Richard and Colleen want to save for a boat. They estimate that they'll need about $4,500 in 7 years. Their bank offers a 2% interest rate compounded semiannually. How much would they need to deposit now in order to have $4,500 in 7 years?
Answer by stanbon(75887)

Richard and Colleen want to save for a boat. They estimate that they'll need about $4,500 in 7 years. Their bank offers a 2% interest rate compounded semiannually. How much would they need to deposit now in order to have $4,500 in 7 years?
--------
A(t) = P(1+(r/n))^(nt)
--------
4500 = P(1+(0.02/2))^(2*7)
-----
4500 = P(1.1)^14
------
P = 4500/1.1^14 = $1184.99
===============================
Cheers,
Stan H.
==========